### v1

# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

Abstract
Model-based Offline Reinforcement Learning trains policies based on offline datasets and model dynamics, without direct real-world environment interactions. However, this method is inherently challenged by distribution shift. Previous approaches have primarily focused on tackling this issue directly leveraging off-policy mechanisms and heuristic uncertainty in model dynamics, but they resulted in inconsistent objectives and lacked a unified theoretical foundation. This paper offers a comprehensive analysis that disentangles the problem into two key components: model bias and policy shift. We provide both theoretical insights and empirical evidence to demonstrate how these factors lead to inaccuracies in value function estimation and impose implicit restrictions on policy learning. To address these challenges, we derive adjustment terms for model bias and policy shift within a unified probabilistic inference framework. These adjustments are seamlessly integrated into the vanilla reward function to create a novel Shifts-aware Reward (SAR), aiming at refining value learning and facilitating policy training. Furthermore, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate the SAR for policy optimization. Empirically, we show that SAR effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks, underscoring its practical effectiveness and validating our theoretical analysis.
footnotetext: footnotetext: Preprint. Under Review.
Introduction
Offline reinforcement learning (RL) (Lange, Gabel, and Riedmiller 2012; Levine et al. 2020) learns policies from a pre-collected dataset generated by a behavior policy, avoiding additional online interaction with the environment. This approach showcases remarkable potential in data-driven decision scenarios (Emerson, Guy, and McConville 2023; Sinha, Mandlekar, and Garg 2022) where the exploration is costly or hazardous. The model-based framework (Luo et al. 2018) is particularly effective for offline RL (Yu et al. 2020, 2021; Sun et al. 2023; Li et al. 2024). It involves learning models of the environment from the given dataset and utilizing these models to generate data for policy training. Nevertheless, directly learning a policy using offline and model-generated data introduces the distribution shift challenge. This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift. For instance, MOPO (Yu et al. 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards. MOBILE (Sun et al. 2023) penalizes value learning of model data according to the uncertainty of off-policy methods in online RL and the Bellman Q-function estimated by ensemble models. RAMBO (Rigter, Lacerda, and Hawes 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions. However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
In this paper, we demystify the distribution shift problem in model-based RL and study its fundamental properties. We first show that this problem arises from both model bias and policy shift, and further analyze their impacts on policy optimization. Unlike the heuristic descriptions of model uncertainty, model bias fundamentally characterizes the inconsistency between the learned model and actual environment dynamics in predicting subsequent states. This bias leads to inaccurate estimations of trajectory returns, resulting in distorted objectives. If the learned model overestimates or underestimates the probability of generating a trajectory, it consequently causes the overestimation or underestimation of trajectory returns. Additionally, policy shift depicts the changes in state-action visitation distribution caused by the deviation between the learned policy $\pi$ and behavior policy $\pi_{b}$. This shift implicitly forces $\pi$ towards $\pi_{b}$, hindering the learning process and slowing down policy convergence. In contrast, previous studies heuristically mitigate these two issues from off-policy algorithms and model uncertainty.
Figure 1: The negative impacts of model bias and policy shift. (a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) to the adjacent grid, with movement obstructed at the boundaries. The initial state is uniformly distributed, and the agent receives rewards upon reaching specific states. (b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias. Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment. (c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift. We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift. The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and preventing its convergence to local optima. Meanwhile, the policy shift modification revises the reward to motivate $\pi$ to escape the limitation of $\pi_{b}$, accelerating policy convergence. Building upon our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL). This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data. Furthermore, it learns a transition classifier and an action classifier to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward. Our contributions include:
  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of value function, while policy shift imposes an implicit policy limitation to hinder policy convergence.
  • We upgrade the vanilla reward into a shifts-aware reward with model bias adjustment and policy shift modification. This shifts-aware reward is used for policy learning to mitigate distribution shift in model-based offline RL.
  • We devise SAMBO-RL, which trains transition and action classifiers to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks.
Related Work
Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin, Yang, and Wang 2021). Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift. For instance, BEAR (Kumar et al. 2019) and BRAC (Wu, Tucker, and Nachum 2019) constrain the learned policy by minimizing different divergences. BCQ (Fujimoto, Meger, and Precup 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy. On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions. CQL (Kumar et al. 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC (Fujimoto and Gu 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto, Hoof, and Meger 2018) objective. Similarly, EDAC (An et al. 2021) and PBRL (Bai et al. 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks. MCQ (Lyu et al. 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al. 2021), achieving high data efficiency (Li et al. 2024). MOPO (Yu et al. 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions. MOReL (Kidambi et al. 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions. COMBO (Yu et al. 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples. RAMBO (Rigter, Lacerda, and Hawes 2022) adopts an adversarial approach, training the policy and model within a robust framework while ensuring accurate transition predictions of the model. CBOP (Jeong et al. 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al. 2018). Lastly, MOBILE (Sun et al. 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble models.
Preliminaries
MDP.
We focus on a Markov decision process (MDP) specified by the tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. Given an MDP, we define the state value function $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
MBPO.
Similar to previous works (Yu et al. 2020, 2021; Rigter, Lacerda, and Hawes 2022; Sun et al. 2023), we focus on model-based policy optimization (MBPO) (Janner et al. 2019), which employs an actor-critic RL algorithm. MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$. Then it optimizes the policy using mini-batches of data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is sampled from the real dataset $\mathcal{D}_{env}$ with probability $f$ and the model dataset $\mathcal{D}_{m}$ with probability $1-f$.
Trajectory Formulation.
To demystify the distribution shift, We consider the trajectory formulation of optimization objectives. Let $p^{\pi}(\cdot)$ represent the probability distribution over trajectories =$\mathcal{H}=\{\tau=(s_{0},a_{0},r_{0},s_{1},...)\}$ generated by executing policy $\pi$ under dynamics $p$. This is given by:
$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ (1)
The training data $\mathcal{D}_{train}$ can be interpreted as being generated by a data collection policy $\pi_{c}$ through rollouts in dynamics $q$. Consequently, the probability of observing trajectory $\tau$ in the training data can be formulated as $q^{\pi_{c}}(\tau)$. The reinforcement learning objective and the practical training objective are then formalized as the following equations:
$\mathcal{J}_{M}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\right]\text{ and }\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim q^{\pi_{c}}}[R(\tau)],$ (2)
where $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ represents the discounted return of the whole trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$.
Shifts-aware Policy Optimization
We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments. To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward. Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shift.
Demystifying Distribution Shift
We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift. Furthermore, we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence. The supplementary analysis is in Appendix A1.
Shift Weighting.
The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}(\tau)}\left[\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}R(\tau)\right].$ (3)
Compared to (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation.
Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following equation:
$\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}=\underbrace{\prod_{t}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}}_{\text{model bias}}\underbrace{\prod_{t}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}}_{\text{policy shift}}.$ (4)
Model bias illustrates the discrepancy between the dynamics used for training and those of the actual environment, while policy shift represents the divergence between the policies used for sampling training data and those being learned.
Negative Impact of Shift Weighting.
We analyze the impact of shift weighting from model bias and policy shift.
(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from actual environment dynamics $p$. In this context, the training objective $\mathcal{J}(\pi)$ is:
$\mathcal{J}(\pi)$ $=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}\right].$
Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective. Specifically, when the model $q$ overestimates or underestimates the likelihood of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment. This essentially introduces a bias for the optimization objective, resulting in wrong training directions.
(ii) To delve deeper into the policy optimization process under policy shift, we consider the setting where a sampling policy $\pi_{c}$ is executed within the true environment transition dynamics $p$. In this scenario, the training objective $\mathcal{J}(\pi)$ is:
$\mathcal{J}(\pi)$ $=\mathbb{E}_{(s,a)\sim d^{p,\pi}}\left[r(s,a)\frac{d^{p,\pi_{c}}(s,a)}{d^{p,\pi}(s,a)}\right].$
Here, $d^{p,\pi}(s,a)$ is the state-action visitation distribution under dynamics $p$ and policy $\pi$, i.e. $d^{p,\pi}(s,a)=(1-\gamma)\mathbb{E}_{s_{0}\sim\mu_{0},s_{t}\sim p,a_{t}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}(s_{t}=s,a_{t}=a)\right]$, where $\mathbb{I}$ represents the indicator function. Note that $\pi_{c}$ is related to $\pi$. Compared to the true objective, the policy shift implicitly forces policy $\pi$ to approach $\pi_{c}$, hindering the learning process and slowing down policy convergence. Specifically, for a state-action pair $(s,a)$ associated with a high reward, ideal policy updates should increase $\pi(a|s)$. However, if the currently learned policy $\pi$ adequately approximates the optimal policy and satisfies $d^{p,\pi}(s,a)>d^{p,\pi_{c}}(s,a)$, the reward for $(s,a)$ can be underestimated. This situation could even prompt policy updates to decrease $\pi(a|s)$, contrary to the optimal direction. Similarly, for the state-action pair $(s,a)$ with a low reward, if the currently $\pi$ is reasonably good and leads to $d^{p,\pi}(s,a)<d^{p,\pi_{c}}(s,a)$, the reward can be overestimated, potentially resulting in policy updates to increase $\pi(a|s)$ that do not align with optimal policy directions.
Toy Examples.
To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example, as depicted in Fig. 1 (a). In Fig. 1 (b), we compare the training process with and without model bias and find the value function overestimation and underestimation under model bias. These results reveal that even slight deviations in model dynamics can significantly distort value function estimates, leading to a biased training objective. Consequently, the policies learned by the agents deviate from the optimal policy. To assess the impact of the policy shift, we examine a model-free offline setting where data is pre-collected using a behavior policy $\pi_{b}$. Fig. 1 (c) showcases that the KL divergence between the learned policy $\pi$ and $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably. This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
These analyses and empirical evidence highlight the critical role of addressing model bias and policy shift in achieving accurate value learning and efficient policy training.
Shifts-aware Reward via Probabilistic Inference
This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective. We further examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
Shifts-aware Reward.
We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories. It comprises the vanilla reward, adjustments for model bias, and modifications for policy shift, expressed as the following:
$\begin{split}&\tilde{r}(s_{t},a_{t},s_{t+1})=\log r(s_{t},a_{t})\\ &+\frac{1}{(1-\gamma)\gamma^{t}}\left[\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right].\end{split}$ (5)
This modified reward forms the basis of our surrogate objective function during training, which is formulated as:
$\mathcal{L}(\pi)=\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}(s_{t},a_{t},s_{t+1})\right].$ (6)
Additionally, we establish that $\mathcal{L}(\pi)$ serves as a lower bound for the true objective, ensuring its rationality as a surrogate.
Theorem 1.
Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
$\log\left[\mathcal{J}_{M}(\pi)\right]\geq(1-\gamma)\mathcal{L}(\pi),\quad\forall\ \pi.$
Optimization objective based on this shifts-aware reward provides a methodological foundation for integrating shift adjustments seamlessly into reinforcement learning algorithms to uniformly address model bias and policy shift.
Modal Bias Adjustment.
We now examine how this adjustment refines the value estimation and addresses the challenge of model bias. By rolling out the learned policy $\pi$ within the model dynamics $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$. Furthermore, the modified training objective of the model dataset $\mathcal{D}_{m}$ can be computed as the following:
$\tilde{J}(\pi)=\mathbb{E}_{s\sim\mathcal{D}_{m},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{m}}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\mathbb{E}_{(s,a)\sim\mathcal{D}_{m}}\left[\log r(s,a)-\alpha D_{\operatorname{KL}}(m(\cdot|s,a)\|p(\cdot|s,a))\right].$
This modified reward consists of the logarithmic vanilla reward penalized by the KL divergence between the model and environment dynamics. This adjustment reduces value function misestimations by accounting for the accuracy of the learned model, encouraging the development of both a high-return policy and an accurate model approximation.
Policy Shift Modification.
To delve into the positive effect of policy shift modification, we consider a dataset $\mathcal{D}_{b}$ generated by executing the behavior policy $\pi_{b}$ within the environment $p$. Moreover, the modified training objective of the dataset $\mathcal{D}_{b}$ is derived as the following formulation:
$\tilde{J}(\pi)=\mathbb{E}_{s\sim\mathcal{D}_{b},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{b}}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\mathbb{E}_{s\sim\mathcal{D}_{b}}\left[\mathbb{E}_{a\sim\pi(\cdot|s)}\left[\log r(s,a)\right]+\beta D_{\operatorname{KL}}(\pi(\cdot|s)||\pi_{b}(\cdot|s))\right],$
which introduces a bonus to the vanilla reward based on the KL divergence between $\pi$ and $\pi_{b}$. This modification guides $\pi$ to break out of the confines of the behavior policy $\pi_{b}$, promoting better exploration and more effective learning.
Shifts-aware Model-based Offline RL
We apply the shifts-aware reward in model-based offline RL and develop the practical framework, SAMBO-RL.
To adapt the shifts-aware reward for practical implementation, where training data is typically stored in transition format rather than trajectories, we introduce two weighting factors, $\alpha$ and $\beta$. These factors replace the respective coefficients in the reward function, shaping it as the following:
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}+\beta\log\frac{\pi(a|s)}{\pi_{c}(a|s)}.$
These weighting factors ensure the reward and adjustments are properly balanced in transition-based learning scenarios, which is empirically validated in the following experiments.
Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al. 2020, 2022). Details are shown in Appendix A3. Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively. The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$.
We train the classifiers by minimizing the loss functions:
$\mathcal{L}(\varphi)=-\mathbb{E}_{\mathcal{D}_{1}}[\log C_{\varphi}]-\mathbb{E}_{\mathcal{D}_{2}}[\log(1-C_{\varphi})].$ (7)
For the transition classifier $C_{\varphi}=C_{\phi}(s,a,s^{\prime})$, $\mathcal{D}_{1}=\mathcal{D}_{env}$ is the offline dataset, and $\mathcal{D}_{2}=\mathcal{D}_{m}$ is the model-generated dataset. For the action classifier $C_{\varphi}=C_{\psi}(s,a)$, $\mathcal{D}_{1}=\mathcal{D}_{\pi}$ is the dataset collected by executing the current policy $\pi$, and $\mathcal{D}_{2}=\mathcal{D}_{env}$ is the offline dataset. Based on these classifiers, we approximate the shifts-aware reward as the following:
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)$ (8)
$+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
The complete procedure of SAMBO-RL is outlined in Algorithm 1, with further details provided in Appendix B.
Algorithm 1 SAMBO-RL
Input: Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, shifts-aware reward hyperparameters $\alpha$ and $\beta$.
1:  Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
2:  Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
3:  for $i=1,2,...,N_{iter}$ do
4:     Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
5:     Generate synthetic rollouts by model $m_{\theta}$. Add transition data in these rollouts to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
6:     Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (7).
7:     Compute shifts-aware reward according to Eq. (8).
8:     Run SAC (Haarnoja et al. 2018) with shifts-aware reward to update policy $\pi$.
9:  end for
Experiments
In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous methods, and (3) analyzing the contribution of each component and hyperparameter within SAMBO to its overall performance.
We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance. Our implementation is based on the OfflineRL-Kit library***https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of our algorithm are consistent with the settings in this library.
Effectiveness of Shifts-aware Reward
(a) Model Bias Adjustment
(b) Policy Shift Modification
Figure 2: Assessment of the shifts-aware reward. (a) Model bias adjustment mitigates inaccuracies in value estimation, achieving a nearly optimal return. (b) Policy shift modification addresses implicit policy restrictions, leading to faster convergence.
We conduct an evaluation of the effectiveness of the Shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid environment illustrated in Fig. 1 (a).
Model Bias Adjustment.
Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective. To evaluate the effectiveness of model bias adjustment, we examine its amendment to the training process. As illustrated in Fig. 2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM), the shifts-aware reward with the overestimating model (SAR with OM), the vanilla reward with the underestimating model (UM), and the shifts-aware reward with the underestimating model (SAR with UM). We then measure the expected return of the learned policy after each update during training in the actual environment. The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy. In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, thereby sustaining optimal returns. Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving near highest return.
Policy Shift Modification.
Policy shift implicitly restricts policy updates. To assess the effectiveness of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig. 1 (a). We test four settings: using either the vanilla reward or the shifts-aware reward, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy. We measure the KL divergence between the learned policy and the behavior policy after each update. As illustrated in Fig. 2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence. In contrast, the algorithm employing the shifts-aware reward effectively motivates policy to be aggressive and optimistic and overcomes this restriction, displaying a higher and more stable divergence and achieving faster policy convergence.
Results on the D4RL Benchmark
Datasets.
In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al. 2020) on the MuJoCo simulator (Todorov, Erez, and Tassa 2012). The evaluation encompasses 12 datasets spanning three environments (HalfCheetah, Hopper, Walker2d) and four dataset types (random, medium, medium-replay, medium-expert). The “v2” versions of datasets are employed for assessment.
Baselines.
We compare SAMBO against multiple baselines, including both model-free and model-based offline RL algorithms. For model-free approaches, BC (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011) learns behavior policy, CQL (Kumar et al. 2020) penalizes the Q-function according to policy change, TD3+BC (Fujimoto and Gu 2021) adds a BC term to regularize the policy, and EDAC (An et al. 2021) quantifies the uncertainty of the Q-value using an ensemble of Q-function networks. For model-based approaches, MOPO (Yu et al. 2020) penalizes rewards via the predicted variance of models. COMBO (Yu et al. 2021) implements CQL within a model-based framework. TT (Janner, Li, and Levine 2021) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter, Lacerda, and Hawes 2022) adversarially trains the policy and the dynamics model within a robust framework. MOBILE (Sun et al. 2023) penalizes the value learning of model data based on the uncertainty of the Bellman Q-function, derived from an ensemble of models.
Comparison Results.
Results are presented in Table 1. Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al. 2020). SAMBO achieves the highest scores on two random datasets and demonstrates strong performance across both medium and medium-replay datasets. Additionally, SAMBO shows competitive results on medium-expert datasets. This performance aligns with our theoretical analysis. By addressing both model bias and policy shift, SAMBO effectively balances the exploitation of offline data with the need to overcome the limitations of the behavioral policy. This approach enables the algorithm to be more aggressive and optimistic in its search for the optimal policy. In scenarios involving low-quality datasets, the aggressive exploration of SAMBO is more likely to yield benefits, potentially leading to the discovery of a superior policy. Conversely, a conservative approach may impede the learning of optimal policy in such contexts. For high-quality datasets, the advantages of aggressive exploration tend to diminish. Fortunately, in these scenarios, methods that emphasize estimating Q-value uncertainty perform effectively. This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO TT RAMBO MOBILE SAMBO (Ours)
halfcheetah-random 2.2 31.3 11.0 28.4 35.4 38.5 38.8 6.1 39.5 39.3 39.7$\pm$2.0
hopper-random 3.7 5.3 8.5 25.3 11.7 31.7 17.9 6.9 25.4 31.9 32.4$\pm$0.5
walker2d-random 1.3 5.4 1.6 16.6 13.6 7.4 7.0 5.9 0.0 17.9 7.0$\pm$8.2
halfcheetah-medium 43.2 46.9 48.3 65.9 42.3 72.4 54.2 46.9 77.9 74.6 72.5$\pm$3.8
hopper-medium 54.1 61.9 59.3 101.6 28.0 62.8 97.2 67.1 87.0 106.6 99.7$\pm$0.3
walker2d-medium 70.9 79.5 83.7 92.5 17.8 84.1 81.9 81.3 84.9 87.7 78.3$\pm$2.6
halfcheetah-medium-replay 37.6 45.3 44.6 61.3 53.1 72.1 55.1 44.1 68.7 71.7 70.6$\pm$1.6
hopper-medium-replay 16.6 86.3 60.9 110.0 67.5 92.7 89.5 99.4 99.5 103.9 101.5$\pm$1.5
walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 82.8$\pm$0.2
halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 94.8$\pm$2.8
hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 106.5$\pm$5.6
walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 80.2$\pm$6.8
Table 1: Results on the D4RL Gym benchmark. MOPO’s results are obtained on the “v0” datasets; MOPO*’s results are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets. The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. The top three scores are highlighted in bold. We use an asterisk “*” for the best score, an underscore “_” for the second, and bold text for the third.
Ablation Studies
We conducted ablation studies on our method to understand why SAMBO demonstrates competitive performance. The following experimental results are presented in Figure 3.
(a) Ablation study for SAMBO
(b) Hyperparameter tuning for $\alpha$
(c) Hyperparameter tuning for $\beta$
Figure 3: Illustration of hyperparameter experiments in the HalfCheetah task with a random dataset. (a) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward. (b) A large model bias adjustment coefficient $\alpha$ increases conservativeness. (c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
Ablation.
Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification. To elucidate the importance of these components, we perform ablation studies on the Halfcheetah task. We assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent. The ablation results are presented in Figure. 3 (a).
We find that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates. This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration. However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance. In contrast, SAMBO incorporates both model bias adjustment and policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance. This underscores the effectiveness of model bias adjustment.
The ablation experiment wo/ps incorporates model bias adjustment but does not address policy shift. As a result, while wo/ps mitigates value misestimation caused by model bias, the lack of policy shift modification causes the algorithm to remain implicitly constrained by the behavioral policy. Therefore, wo/ps demonstrates reduced effectiveness in discovering a more optimal policy. In contrast, SAMBO integrates policy shift modification during conservative updates, which enables more aggressive exploration and results in superior performance compared to wo/ps. This highlights the effectiveness of incorporating policy shift modification.
Hyperparameter Tuning.
SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
The model bias adjustment coefficient $\alpha$ regulates the extent of adjustment applied to model-generated data. A larger $\alpha$ imposes a significant penalty on the model, leading to excessively conservative utilization of the model. Consequently, even in regions where the model performs well, exploration may be restricted due to this excessive penalty, potentially resulting in suboptimal performance. Conversely, a smaller $\alpha$ can cause the algorithm to be overly optimistic about exploring risky regions, thereby impairing performance. To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task. The results, shown in Fig. 3 (b), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance. Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration. This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy. A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy. However, this also introduces a heightened risk of taking unsafe actions. As illustrated in Fig. 3 (c), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness. A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability. Therefore, $\beta$ also needs to be set appropriately to balance exploration and stability.
Conclusion and Discussion
In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift. Model bias results in distorted objectives, while policy shift obstructs the convergence of policies. We introduce a shifts-aware policy optimization framework to uniformly address these challenges and apply it within the model-based offline setting. Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms. Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training. In future work, We aim to explore alternative methods that directly address distribution shift using datasets, eliminating the need for classifier estimation. Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems. Our novel analysis and results related to policy shift in this context open new avenues for better exploiting offline datasets and overcoming the restrictions imposed by behavior policies.
Appendix A A. Theoretical Analysis
A1. The Supplement Analysis of Distribution Shift
Dual Formulation of Model Bias.
The reinforcement learning optimization objective can be equivalently reformulated in its dual form, expressed as $\mathcal{J}_{M}(\pi)=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)]$, where $d^{p,\pi}(s,a)$ denotes the state-action visitation distribution under dynamics $p$ and policy $\pi$. To illustrate the impact of model bias in the dual formulation, we similarly examine the scenario where the training policy $\pi$ is applied within the model dynamics $q$. The training objective $\mathcal{J}(\pi)$ is derived as follows:
$\mathcal{J}(\pi)=\mathbb{E}_{(s,a)\sim d^{q,\pi}}[r(s,a)]=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)\frac{d^{q,\pi}(s,a)}{d^{p,\pi}(s,a)}].$
Model bias can result in inaccuracies in estimating the state-action visitation distribution, subsequently affecting the accuracy of value function estimation. Similarly, if the learned model overestimates or underestimates the state-action visitation distribution, it will lead to corresponding overestimations or underestimations of rewards, thereby distorting the objective.
Trajectory Formulation of Policy Shift.
Similarly, we consider the trajectory formulation of policy shift. We execute the sampling policy $\pi_{c}$ within the environment transition dynamics $p$. Then the training objective $\mathcal{J}(\pi)$ is then given by:
$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi_{c}}}[R(\tau)]=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}\right].$
The policy shift can influence the value estimation of trajectories, imposing implicit constraints on policy learning. For a high-reward trajectory $\tau$, if the current policy $\pi$ closely approximates the optimal policy and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})>\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, it may lead to an underestimation of the return for trajectory $\tau$. Conversely, for a low-reward trajectory $\tau$, if the current policy is relatively effective and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})<\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, this can result in an overestimation of the return for trajectory $\tau$. Both of these impede the effective learning of the policy.
A2. The Proof of Theorem 1
Theorem 1.
Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
$\log\left[\mathcal{J}_{M}(\pi)\right]\geq(1-\gamma)\mathcal{L}(\pi),\quad\forall\ \pi.$
Proof.
The cumulative discounted reward $R(\tau)$ of trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$ can be written as:
$R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})=\frac{1}{1-\gamma}\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}r(s_{t},a_{t}).$ (9)
As discussed in the main text, the probability distribution over a trajectory $\tau$ can be formulated as follows:
$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})\text{ , }q^{\pi_{c}}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t}).$ (10)
Then according to (9) and (10) and Jensen’s inequality, we have
$\log\left[\mathcal{J}_{M}(\pi)\right]$ $=\log\mathbb{E}_{p^{\pi}(\tau)}R(\tau)$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{p^{\pi}(\tau)}{q^{\pi_{c}}(\tau)}\right]$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{\mu(s_{0})\prod_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})}{\mu(s_{0})\prod_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t})}\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log R(\tau)+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log\frac{1}{1-\gamma}+\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\left((1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$=(1-\gamma)\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\log r(s_{t},a_{t})+\frac{1}{(1-\gamma)\gamma^{t}}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right)\right]$
$=(1-\gamma)\mathcal{L}(\pi).$
A3. Classifiers Estimation Derivations
We assume that all steps in the derivation are appropriate, that the denominator is non-zero, and that the argument of the logarithm is not zero.
Transition Classifier Derivation.
Define two events $\mathbb{E}$ and $\mathbb{M}$, where
[…]
The transition classifier is defined as $C_{\phi}(s_{t},a_{t},s_{t+1})=P(\mathbb{E}|s_{t},a_{t},s_{t+1})$. Then
$C_{\phi}(s,a,s_{t+1})=P(\mathbb{E}|s_{t},a_{t},s_{t+1})=$ $\frac{P(\mathbb{E},s_{t},a_{t})P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t},a_{t},s_{t+1})}$
$=\frac{P(\mathbb{E}|s_{t},a_{t})}{P(s_{t},a_{t})}\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t},a_{t},s_{t+1})}.$ (11)
Events $\mathbb{E}$ and $\mathbb{M}$ form a partition of the probability space, which means that
$P(\mathbb{M}|s_{t},a_{t},s_{t+1})=1-P(\mathbb{E}|s_{t},a_{t},s_{t+1})=1-C_{\phi}(s,a,s_{t+1}).$
Similarly, we have
$1-C_{\phi}(s,a,s_{t+1})$ $=\frac{P(\mathbb{M},s_{t},a_{t})P(s_{t+1}|s_{t},a_{t},\mathbb{M})}{P(s_{t},a_{t},s_{t+1})}$
$=\frac{P(\mathbb{M}|s_{t},a_{t})}{P(s_{t},a_{t})}\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{M})}{P(s_{t},a_{t},s_{t+1})}.$ (12)
According to (11) and (12), we have
$\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t+1}|s_{t},a_{t},\mathbb{M})}\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (13)
Specifically, $P(s_{t+1}|s_{t},a_{t},\mathbb{E})=p(s_{t+1}|s_{t},a_{t})$, where $p$ represents the environment transition dynamic; $P(s_{t+1}|s_{t},a_{t},\mathbb{M})=q(s_{t+1}|s_{t},a_{t})$, where $q$ denotes the model transition dynamic. Then taking the logarithm of both sides of (13), we obtain:
$\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (14)
Given the independence of the state-action pair $(s_{t},a_{t})$ from the transition dynamics, we assume that the visit distribution of $(s_{t},a_{t})$ remains identical. Then
$\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}\approx\frac{d^{\pi}(s_{t},a_{t})\left|\mathcal{D}_{env}\right|}{d^{\pi}(s_{t},a_{t})\left|\mathcal{D}_{m}\right|}=\frac{\left|\mathcal{D}_{env}\right|}{\left|\mathcal{D}_{m}\right|},$ (15)
where $\left|\mathcal{D}_{env}\right|$ and $\left|\mathcal{D}_{m}\right|$ represent the sizes of the datasets $\mathcal{D}_{env}$ and $\mathcal{D}_{m}$, respectively. Thus, $\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}$ can be approximated by a constant. Then according to (14) and (15), we have
$\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}\approx\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}-c_{1},$
where $c_{1}$ is a constant that does not influence the training of SAMBO.
Action Classifier Derivation.
The derivation of the action classifier is similar to that of the transition classifier. Define two events $\mathbb{A}$ and $\mathbb{C}$, where
[…]
={Given the state $\mathbb{C}=\left\{\text{Given the state }s_{t},\text{ the agent takes action }a_{t}\text{ according to the behavior policy }\pi_{c}\right\}.$
The action classifier is defined as $C_{\psi}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
$C_{\psi}(s,a)=P(\mathbb{A}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{A})P(\mathbb{A}|s_{t})P(s_{t})}{P(s_{t},a_{t})},$ (16)
$1-C_{\psi}(s,a)=P(\mathbb{C}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{C})P(\mathbb{C}|s_{t})P(s_{t})}{P(s_{t},a_{t})}.$ (17)
Specifically, $P(a_{t}|s_{t},\mathbb{A})=\pi(a_{t}|s_{t})$, $P(a_{t}|s_{t},\mathbb{C})=\pi_{c}(a_{t}|s_{t})$. Then according to (16) and (17), we have
$\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}=\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}.$ (18)
Similarly, $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}$ can be also approximated by the constant $\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|}$. According to the sampling method in Algorithm 2, we have
$\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}=\frac{d(s_{t})\left|\mathcal{D}_{\pi}\right|}{d(s_{t})\left|\mathcal{D}_{env}\right|}=\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|},$
where $d(s_{t})$ denotes the distribution of states in the offline dataset, and $\mathcal{D}_{\pi}$ represents the dataset collected by policy $\pi$. Thus, taking the logarithm of both sides of (18), we get
$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}-c_{2},$
where $c_{2}$ is a constant that does not influence the training process.
Appendix B B. Implementation Details
We present a detailed outline of the SAMBO algorithm in Algorithm 2.
Algorithm 2 SAMBO-RL
Input: Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, shifts-aware reward hyperparameters $\alpha$ and $\beta$, rollout horizon $h$, rollout batch size $b$.
1:  Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
2:  Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
3:  for $i=1,2,...,N_{iter}$ do
4:     Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
5:     for $1,2,...,b$ (in parallel) do
6:        Sample state $s_{1}$ from $\mathcal{D}_{env}$ for the initialization of the rollout.
7:        for $j=1,2,...,h$ do
8:           Sample an action $a_{j}\sim\pi(s_{j})$.
9:           Randomly pick dynamics $m$ from $\{m_{\theta}^{i}\}_{i=1}^{N}$ and sample $s_{j+1},r_{j}\sim m(s_{j},a_{j})$.
10:           Add sample $(s_{j},a_{j},r_{j},s_{j+1})$ to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
11:        end for
12:     end for
13:     Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (19) and Eq. (20), respectively.
14:     Drawing samples from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, compute shifts-aware reward according to Eq. (21) and Eq. (22), respectively.
15:     Run SAC with shifts-aware reward to update policy $\pi$.
16:  end for
B1. Model Training
In our approach, the model is represented by a probabilistic neural network that outputs a Gaussian distribution for both the next state and reward, given the current state and action:
$m_{\theta}\left(s_{t+1},r_{t}\mid s_{t},a_{t}\right)=\mathcal{N}\left(\mu_{\theta}\left(s_{t},a_{t}\right),\Sigma_{\theta}\left(s_{t},a_{t}\right)\right).$
Our model training approach is consistent with the methodology used in prior works (Janner et al. 2019; Yu et al. 2020). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. During model rollouts, we randomly choose one model from the best five models.
B2. Classifiers Training
The transition classifier $C_{\phi}(s,a,s^{\prime})\in[0,1]$ represents the probability that the transition originated from the environment. It is trained by minimizing the standard cross-entropy loss function:
$\mathcal{L}_{\phi}=-\mathbb{E}_{\mathcal{D}_{env}}\left[\log C_{\phi}(s,a,s^{\prime})\right]-\mathbb{E}_{\mathcal{D}_{m}}\left[\log(1-C_{\phi}(s,a,s^{\prime}))\right].$ (19)
The action classifier $C_{\psi}(s,a)\in[0,1]$ represents the probability that a given state-action pair $(s,a)$ was generated by the current policy $\pi$. This classifier is trained by minimizing:
$\mathcal{L}_{\psi}=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi}(s,a)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi}(s,a))\right].$ (20)
For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $q$. Furthermore, we approximate the training data collection policy for $\mathcal{D}_{m}$ using the current policy $\pi$, i.e. $\pi_{c}=\pi$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{m}$, the shifts-aware reward simplifies to
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}.$
Then, for all $(s,a,s^{\prime})\in\mathcal{D}_{m}$, the shift-aware reward can be estimated using the following formulation,
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}.$ (21)
For data in the offline dataset $\mathcal{D}_{env}$, the dynamics are governed by environment $p$, i.e. $q=p$, and the training data collection policy aligns with the behavioral policy, i.e. $\pi_{c}=\pi_{b}$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{env}$, the shifts-aware reward simplifies to
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\beta\log\frac{\pi(a|s)}{\pi_{b}(a|s)},$
which is independent of the subsequent state $s^{\prime}$. Then, for all $(s,a,s^{\prime})\in\mathcal{D}_{env}$, the shift-aware reward can be estimated by
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$ (22)
B3. Policy Learning
The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al. 2023). For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$. The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 2.
Hyperparameters Value Description
K 2 The number of critics.
$\gamma$ 0.99 Discount factor.
$l_{r}$ of actor $1\times 10^{-4}$ Policy learning rate.
$l_{r}$ of critic $3\times 10^{-4}$ Critic learning rate.
Optimizer Adam Optimizers of the actor and critics.
$f$ 0.05 Ratio of the real samples.
Batch size 256 Batch size for each update.
$N_{\text{iter}}$ 3M Total gradient steps.
Table 2: Hyperparameters for policy optimization in SAMBO
Appendix C C. Experimental Details
C1. Benchmark
We conduct experiments on Gym tasks (“v2” version) within the D4RL benchmark and now provide the sources for the reported performance on this benchmark. For TD3+BC (Fujimoto and Gu 2021), EDAC (An et al. 2021), TT (Janner, Li, and Levine 2021), RAMBO (Rigter, Lacerda, and Hawes 2022), and MOBILE (Sun et al. 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym. Since COMBO (Yu et al. 2021) does not provide source code, we include the results as reported in the original paper. The results of CQL (Kumar et al. 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym. The results of MOPO* are obtained from experiments using the OfflineRL-Kit libraryhttps://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets. These scores are referenced in Table 1.
C2. Hyperparameters
We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 3.
Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
halfcheetah-random 0.01 0.01 5 -0.2
hopper-random 0.02 0.01 5 0
walker2d-random 0.03 0.03 5 0
halfcheetah-medium 0.01 0.01 5 0
hopper-medium 0.02 0.01 5 0
walker2d-medium 0.035 0.01 5 0
halfcheetah-medium-replay 0.01 0.01 5 0
hopper-medium-replay 0.02 0.01 5 0
walker2d-medium-replay 0.01 0.01 1 -0.2
halfcheetah-medium-expert 0.04 0.01 5 0
hopper-medium-expert 0.04 0.02 5 0
walker2d-medium-expert 0.05 0.05 1 0
Table 3: Hyperparameters of SAMBO used in the D4RL datasets.
Model Bias Coefficient $\alpha$.
We tune $\alpha$ in the range of $[0.01,0.05]$. For training stability, in the walker2d-medium dataset, we fix the transition classifier after 1 million iterations by setting $\alpha=0$. In the walker2d-medium-expert dataset, we apply an annealing factor to the transition classifier. Initially set to 3, the factor remains constant for the first 1 million iterations and then decays reciprocally to 1 over the subsequent 2 million iterations.
Policy Shift Coefficient $\beta$.
We tune $\beta$ in the range of $[0.01,0.05]$.
Rollout Length $h$.
Similar to MOPO, We perform short-horizon branch rollouts in SAMBO. we tune $h$ in the range of $\{1,5\}$ for Gym tasks in the D4RL benchmark.
Reward Truncation Coefficient $c$.
To compute the logarithm of the reward term in the shifts-aware reward, we apply a translation to the real reward, which does not affect policy optimization. Specifically, given the offline dataset $\mathcal{D}_{env}$, let $r_{max}$ denote the maximum reward and $r_{min}$ denote the minimum reward in this dataset. Then we perform a translation and truncation on the real reward as follows:
$r^{\prime}=\min\{1\times 10^{-8},r-c(r_{max}-r_{min})+1\times 10^{-8}\}.$
Then we calculate the shifts-aware reward by
$\tilde{r}(s,a,s^{\prime})\approx\log r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
In most datasets, we set $c=0$. However, we observe that employing a negative value for $c$ yields favorable results in certain datasets.
C3. Computing Infrastructure
All experiments are conducted using a single GeForce GTX 3090 GPU and an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz. Our implementation code is based on the OfflineRL-Kit library. The software libraries and frameworks utilized in SAMBO are consistent with those specified in this library.